OpenStudy (vortish):

simplify the expression if possible log o.oo1

Directrix (directrix):

Is this supposed to be the base ten log? --> log(.001) ?

Directrix (directrix):

If so, it will simplify. @vortish

Directrix (directrix):

log(.001) = log (1* 10^(-3) ) = log (1) + log ((10^(-3) ) Sorry but I cannot help beyond this point unless you respond with some sort of post about the problem and its solution. @vortish

OpenStudy (anonymous):

Hello? Anyone there?

OpenStudy (vortish):

its just says simplify and log 0.001 =

Directrix (directrix):

Pick up from here: log (1* 10^(-3) ) = log (1) + log ((10^(-3) )

OpenStudy (vortish):

ok i understand that sorry got called to dinner